OpenStudy (anonymous):

series that goes 5/3 + 1/3 + 1/7+..... i need to find the nth term?

OpenStudy (anonymous):

first you need to find out if this is an arithmetic sequence or a geometric sequence

OpenStudy (anonymous):

depending on the type of sequence will let you know what formula to use

OpenStudy (anonymous):

no it's not one of those things, i know those formulas. we're doing this thing where we need to find the rule for n and the sum

OpenStudy (anonymous):

so i guess my question was stated incorrectly, sorry

OpenStudy (anonymous):

that's ok so the formula that I know... and I'm not sure if I've learned this yet, because we just started learning this in my class, but here is my best educated guess of the formula you'd use.\[a _{1}(r)^{i-1}+\sum_{i-1}^{\infty}a _{2}(r)^{i-1}\] where r is the ratio.

OpenStudy (anonymous):

the nth term would be in the a sub 1 and a sub 2. Sorry I'm not that helpful :/

OpenStudy (anonymous):

thanks but i really don't know how to do this still :/ i think the problem is that i don't know what r is, which is the problem

OpenStudy (anonymous):

what i mean is, it's like 1/2 + 1/6 + 1/12. then you see that it's 1/(1*2) + 1/(2*3) + 1/(3*4) so the n thing is 1/n(n+1)

OpenStudy (anonymous):

this looks to be a 'finite geometric series'

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

uhh maybe? i think that might work?

OpenStudy (anonymous):

actually, what are you suppose to do, find the sum of all the terms or to find the equation that gives you the value of the nth term?

OpenStudy (anonymous):

both. but if i have the nth then i can find the sum on my own

OpenStudy (anonymous):

look in your book for 'finite geometric series' it will as you to do a check of 'if r<1' if so, then the series converges. I don't have my book, and don't trust the web, so just find that equation and enter in the info the problem gave you to find the sum of the series

OpenStudy (anonymous):

\[S _{n} = \frac{ a _{1}(1-r ^{n}) }{ 1-r }\]

OpenStudy (anonymous):

and the info the problem gave me is those 3 numbers. that's all i have.

OpenStudy (anonymous):

so thats the finite equation? Not the infinate?

OpenStudy (anonymous):

r = a2/a1